### v1

@@ Abstract
The accurate modeling of semiconductor devices plays a critical role in the development of new technology nodes and next-generation devices. Semiconductor device designers largely rely on advanced simulation software to solve the drift-diffusion equations, a coupled system of nonlinear partial differential equations that describe carrier transport in semiconductor devices. While these tools perform well for forward modeling, they are not suitable to address inverse problems, for example, determining doping profiles, material, and geometrical parameters given a desired device performance. Meanwhile, physics-informed neural networks (PINNs) have grown in popularity in recent years thanks to their ability to efficiently and accurately solve inverse problems at minimal computational cost compared to forward problems. In this study, we introduce the Drift-Diffusion Network (DDNet), a unified physics-informed deep learning solver for the forward and inverse mesh-free solutions of the drift-diffusion equations of semiconductor device modeling. Using prototypical device configurations in one- and two spatial dimensions, we show that DDNet achieves low absolute and relative error compared to traditional simulation software while additionally solving user-defined inverse problems with minimal computational overhead. We expect that DDNet will benefit semiconductor device modeling by facilitating exploration and discovery of novel device structures across comprehensive parameter sets in a fully automated way.
@@ 1 Introduction
Semiconductor devices are fundamental components of modern microelectronics, enabling the operation of integrated circuits, optoelectronic systems, and power management technologies [1]. Their behavior, governed by charge carrier transport at the microscale [2], directly impacts the performance, energy efficiency, and scalability of computing, communication, and sensing technologies [3, 4, 5].
@@
As device architectures continue to scale and diversify, precise modeling of semiconductor physics remains essential for both scientific understanding and technological development. Among the various approaches, the drift-diffusion model [6, 7, 8], consisting of a set of coupled parabolic partial differential equations (PDEs), forms the backbone of most modern commercial simulation tools. These equations, typically grouped into Poisson and continuity equations, describe the response of charge carriers under external excitation. Due to their intrinsic nonlinearity, analytical solutions are only possible in limited cases, and numerical approaches are required [9].
@@
Commercial software packages such as Synopsys TCAD Sentaurus [10], Silvaco Atlas [11], and COMSOL Multiphysics [12] implement advanced numerical solvers based on finite element or finite volume methods. These tools provide high accuracy in 1D, 2D and 3D simulations, support a wide range of physical models and boundary conditions, and are widely used for device design, optimization, and analysis in both academia and industry.
@@
Despite their maturity, traditional solvers face several limitations. Their convergence often depends on mesh quality, numerical approximations of differential operators, initial guesses, and extensive parameter tuning. Furthermore, solving coupled nonlinear PDEs typically requires Newton-type second-order methods, which are highly sensitive to initialization choices [9]. Lastly, exploring high-dimensional parameter spaces, solving inverse problems, or integrating sparse experimental data often exceeds the capabilities of conventional solvers. These challenges motivate the development of an alternative paradigm that takes advantage of the flexibility and power of artificial neural networks (ANNs) and provides physical consistency, numerical accuracy, scalability, parameter training capabilities, as well as the efficient solution of user-defined inverse problems [13].
@@
In this context, recent advances in scientific machine learning have introduced physics-informed neural networks (PINNs) as a mesh-free alternative for solving forward and inverse problems governed by PDEs [13, 14, 15, 16, 17]. Unlike conventional data-driven neural networks, PINNs enforce the validity of physical laws by embedding the governing equations directly into the loss function. These models use feedforward neural network architectures as trainable surrogates of the PDE solutions and leverage automatic differentiation (AD) to compute the PDE residuals at randomly distributed collocation points in the solution domain and on its boundary [18, 19, 20].
@@
In the case of forward simulations, PINNs eliminate the need for labeled data, relying solely on physics-based constraints. This makes them well-suited for problems where data are scarce or unavailable. Furthermore, inverse problems [21, 22] can be handled by simply introducing an additional loss function, without significant computational overhead [23, 14, 16]. These characteristics make PINNs a flexible and powerful simulations framework for solving a wide range of forward and inverse differential and integro-differential problems.
@@
With respect to semiconductor device modeling, limited progress has been achieved so far using PINN methods [24]. Cao et al. investigated PINNs for Poisson-like equations, without however extending their results to relevant electrical engineering problems [25]. Cai et al. proposed a TCAD forward simulation PINN-based architecture for solving a one-dimensional $p$-$n$ junction simulation. However, they did not discuss the generality of their method in higher dimensions or for inverse problems [26]. Kim and Shin used a PINN-DeepONet hybrid solver to simulate the electrical characteristics of a one-dimensional device starting from from TCAD training data [27]. Lastly, Liu et al. developed a one-dimensional Poisson-Boltzmann solver, which addressed parameter retrieval problems [28]. As a result, there remains a critical need for a forward and inverse (unified) PINN framework that can address relevant devices and coupled simulations with inverse design capabilities. Such a framework should match the precision and reliability of state-of-the-art industrial solvers while demonstrating greater flexibility and scalability in integrating different boundary conditions, generation-recombination rates, and inversely retrieving relevant device parameters. Moreover, an innovative physics-informed deep learning framework should also demonstrate a clear computational advantage in solving relevant nonlinear inverse problems beyond the reach of current numerical methods [29]. As recently pointed out, PINNs should not be employed to solve relatively simple forward problems against a weak baseline without demonstrating a clear computational advantage over traditional solvers [30]. Specifically, we argue that the crucial advantage of PINNs compared to traditional solvers stems from their inherent ability to successfully tackle inverse design problems where physics-constrained solutions, representing targeted functionalities or device configurations, are efficiently recovered even in the case of stiff and ill-posed problems [31, 23]. Furthermore, there have been increasing efforts to extend the applications of physics-based surrogate models to benchmark performance applications, such as in the automotive aerodynamics field [32], stressing the importance of physics-informed deep learning for industrial applications.
@@
In addition, fields such as fluid dynamics and electromagnetics have experienced substantial advances in physics-informed surrogate modeling over the past five years, despite the availability of established forward Navier-Stokes or electrodynamics numerical solvers [12, 14, 15, 16, 17, 23]. In contrast, the comparatively limited progress of PINNs in semiconductor device modeling originates from the intrinsic challenges associated with the highly nonlinear nature of the drift–diffusion equations, whose solutions present large gradients and span extremely large dynamic ranges. In our view, these considerations underscore the need for fundamentally new network architectures and training methodologies.
@@
While DDNet meets the essential benchmarks for comparison with current state-of-the-art modeling tools, it is not intended as a tout court replacement of established numerical methods. Instead, our comparisons should be regarded as seminal steps toward the development of PINN-driven semiconductor device modeling techniques enabling device discovery at minimal computational costs. Ultimately, our results highlight the potential of DDNet as a flexible and generalizable framework for microelectronics research, inverse design, and device discovery, especially in data-scarce and high-dimensional scenarios.
@@ 2 Results
We present here prototypical examples of semiconductor device configurations in order to demonstrate the capabilities of the developed DDNet framework and we compare its predictions with those obtained from state-of-the-art commercial solvers. The following results comprise both forward and inverse design problems, including simulations of 1D and 2D homogeneous silicon junctions under varying bias conditions.
For a device with permittivity $\epsilon$, DDNet solves the general stationary drift-diffusion system for $\phi,n,$ and $p$:
@@
The Poisson-drift-diffusion equations describe charge transport in semiconductors by coupling the Poisson equation for the electrostatic potential $\phi$ to the continuity equations for the electrons ($n$) and holes ($p$) concentrations. To solve Eq. 1 we must specify the doping profile $C=N_{d}^{+}-N_{a}^{-}$, which is the distribution of impurities (dopants) within a semiconductor material, and that determines the electrical properties of the device. $\boldsymbol{J}_{n},\boldsymbol{J}_{p}$ denote the electron and hole current densities, respectively. Their divergences are linked to the generation–recombination term $R$ that represents the net rate of electron–hole pair creation or annihilation per unit volume. In the simulations that follow, we include the Shockley-Read-Hall recombination term, but DDNet has the flexibility to incorporate other mechanisms such as Auger or radiative effects. Charge transport is dominated by two phenomena. Charges driven by the electric field give rise to the drift currents while gradients in charge density originate diffusion currents, with mobilities and diffusivities linked via the Einstein relation [6, 1]. Solving Eq. 1 requires the specification of boundary conditions, which can be Dirichlet, Neumann, or mixed.
2.1 The DDNet framework
DDNet is designed to simultaneously learn the scaled functions $\hat{\phi}$, $\hat{n}$, and $\hat{p}$ [6]. As illustrated in Fig. 1, the architecture comprises three coupled fully-connected deep neural networks (FCNN), namely the $\phi$-Net, $n$-Net, and $p$-Net. In this work, we introduce a logarithmic FCNN architecture for the $n$-Net and $p$-Net subnetworks, which enables the accurate approximation of the large dynamic ranges of the solutions, often spanning twelve to eighteen orders of magnitude. The networks are trained on randomly sampled collocation points in the simulation domain $\Omega$ and its boundary $\partial\Omega$, where the PDE residuals, boundary conditions, and (if applicable) inverse constraints are evaluated and minimized via a cumulative mean-squared error loss. For all the examples that follow, the loss functions of DDNet achieve values lower than $10^{-4}$, without the need for device- or geometry-specific fine-tuning, as shown in the Supplementary Information.
@@ 2.2 One-dimensional $p$-$n$ junction
We first simulate a silicon 1D $p$-$n$ junction at equilibrium using an abrupt doping profile. Despite having an analytical solution [6, 1], this case is computationally challenging due to the presence of a discontinuity in $C$. As shown in Fig. 2a, the junction forms at the interface between $n$-type and $p$-type regions. At thermal equilibrium, the distribution of carriers creates a depletion region and establishes the built-in potential $\phi_{bi}$. Dirichlet Ohmic boundary conditions are applied for all fields to ensure uniqueness of the solution. Fig. 2b compares the DDNet prediction of $\phi$ with TCAD Sentaurus results, showing excellent agreement in logarithmic scale. The carrier density profiles are shown in Fig. 2c–d, achieving relative $L_{1}$ errors below 2%. We also simulated the same junction under forward and reverse bias, and included the results in the Supplementary Information. DDNet accurately captures the evolution of potential and carrier profiles for various applied voltages. In particular, DDNet captures the expected shift of the depletion region. Additionally, we simulate and include in the Supplementary Information an asymmetric ($p^{+}$-$n$) junction, where the depletion region shifts toward the less-doped side.
@@ 2.3 Two-dimensional $p$-$n$ junction
To demonstrate scalability to higher spatial dimensions, we simulate a 2D $p$-$n$ junction with the geometry shown in Fig. 3a and doping profile in Fig. 3b. The boundary conditions are schematized in panel c, where Dirichlet conditions are enforced on source/drain contacts and Neumann conditions elsewhere. Here, the Neumann conditions capture the insulating region, with zero current outflow. The DDNet predictions for $\phi$, $n$, and $p$ are shown in panels e, h, and k, respectively, and compared to TCAD Sentaurus results in panels d, g, and j. The 3D surface plots (panels f, i, l) confirm accurate reconstructions, with the average relative $L_{1}$ errors with TCAD Sentaurus’s solutions falling below 5%. This example demonstrates robust scalability to higher dimensions, highlighting the expressivity of DDNet. In fact, the architecture, collocation points, and loss functions were left unchanged for the solution of the 2D problem, as discussed in the Supplementary Information.
@@
Figure 4: Inverse design capabilities of DDNet. a, Modified DDNet architecture to tackle inverse design and parametric learning tasks. The input space has been expanded to accept device parameters $P_{1},\dots,P_{n}$, and the output space of the $p$-Net and $n$-Net now enables the retrieval of the doping profile $\hat{C}=\hat{N}_{d}^{+}+\hat{N}_{a}^{-}$. b, Example of parametric learning for a pinpin junction, where DDNet extrapolates the maximal electric field behavior as a function of the intrinsically doped region width after training on 4 widths. c, Example of inverse design of a doping profile to achieve the maximal user-defined target performance from only three data points. d, Example of device discovery, where a pnppnp BJT-like structure is retrieved to match the user’s performance criteria.
2.4 Inverse design and parametric learning

### v2

@@ Abstract
The accurate modeling of semiconductor devices plays a critical role in the development of new technology nodes and next-generation devices. Semiconductor device designers largely rely on advanced simulation software to solve drift-diffusion equations, a coupled system of non-linear partial differential equations that describe carrier transport in semiconductor devices. Although these tools perform well for forward modeling, they are not suitable for addressing inverse problems, for example, determining doping profiles, material properties, and geometric parameters given desired device performance. Meanwhile, physics-informed neural networks (PINNs) have grown in popularity in recent years due to their ability to efficiently solve inverse problems at minimal computational cost compared to forward problems. In this study, we introduce the Drift-Diffusion Network (DDNet), a unified physics-informed deep learning solver for forward and inverse mesh-free solutions of the drift-diffusion equations. Using prototypical device configurations in one and two spatial dimensions, we show that DDNet achieves low error compared to traditional simulation software, while additionally solving user-defined inverse problems with minimal computational overhead. We expect that DDNet will benefit semiconductor device modeling by enabling the inverse design paradigm of semiconductor devices from any user-defined objectives, such as electric fields or charge distributions, thus facilitating the discovery of novel structures in comprehensive parameter sets and in a fully automated way.
@@ 1 Introduction
Semiconductor devices are fundamental components of modern microelectronics, enabling the operation of integrated circuits, optoelectronic systems, and power management technologies [sze_physics_2021]. Their behavior, governed by charge carrier transport at the microscale [2016Ferry], directly impacts the performance, energy efficiency, and scalability of computing, communication, and sensing technologies [xie_designing_2024, zhao_review_2024, thennarasu_comprehensive_2025].
@@
As device architectures continue to scale and diversify, precise modeling of semiconductor physics remains essential for both scientific understanding and technological development. Among the various approaches, the drift-diffusion model [1985Markowich, 2012Markowich, 2008Vasileska], consisting of a set of coupled parabolic partial differential equations (PDEs), forms the backbone of most modern commercial simulation tools. These transport equations, typically grouped into Poisson and continuity equations, describe the transport of charge carriers under external excitation. Due to the intrinsic non-linearity of the equations, analytical solutions are only possible in limited cases, and numerical approaches are required [1983BankRose, 1985Markowich, 2012Markowich].
@@
Commercial software packages such as Synopsys TCAD Sentaurus [Sentaurus_Device_W-2024.09-SP1], Silvaco Atlas [silvaco_sftw], and COMSOL Multiphysics [comsol_sftw] implement advanced numerical solvers based on finite element or finite volume methods. These tools provide high accuracy in 1D, 2D and 3D simulations, support a wide range of physical models and boundary conditions, and are widely used for device design, optimization, and analysis in both academia and industry.
@@
Despite their maturity, traditional solvers face several limitations. Their convergence often depends on the quality of the mesh, numerical approximations of differential operators, initial guesses, and extensive parameter tuning. Furthermore, solving coupled non-linear PDEs typically requires Newton-type second-order methods, which are highly sensitive to initialization choices [1983BankRose]. Lastly, exploring high-dimensional parameter spaces, solving inverse problems, or integrating sparse experimental data often exceeds the capabilities of conventional solvers. These challenges motivate the development of an alternative paradigm that takes advantage of the flexibility and power of artificial neural networks (ANNs) and provides physical consistency, scalability, parameter training capabilities, and the fast solution of user-defined inverse design problems [raissi_physics-informed_2019].
@@
In this context, recent advances in scientific machine learning have introduced physics-informed neural networks (PINNs) as a mesh-free alternative to solve forward and inverse problems governed by PDEs [raissi_physics-informed_2019, chen_physics-informed_2020, eivazi_physics-informed_2022, chen_physics-informed_2022, riganti_auxiliary_2023]. Unlike conventional data-driven neural networks, PINNs enforce the validity of physical laws by embedding the governing equations directly into the loss function. These models use feedforward neural network architectures as trainable surrogates of PDE solutions and leverage automatic differentiation (AD) to compute PDE residuals at randomly distributed collocation points in the solution domain and on its boundary [haykin_neural_2009, goodfellow_deep_2016, lu_deepxde_2021].
@@
In the case of forward simulations, PINNs eliminate the need for labeled data, relying solely on physics-based constraints. This makes them well-suited for problems where data are scarce or unavailable. Furthermore, inverse problems [tarantola_inverse_2005, chen_computational_2018] can be handled by simply introducing an additional loss function, without significant computational overhead [riganti_multiscale_2025, chen_physics-informed_2020, chen_physics-informed_2022]. These characteristics make PINNs a flexible and powerful simulation framework for solving a wide range of forward and inverse differential and integro-differential problems [raissi_physics-informed_2019, chen_physics-informed_2020, eivazi_physics-informed_2022, chen_physics-informed_2022, riganti_auxiliary_2023, mao_physics-informed_2020, raissi_hidden_2020, yu_gradient-enhanced_2022, xu_understanding_2025].
@@
With respect to semiconductor device modeling, limited progress has been achieved so far using PINN methods [li_overview_2024]. Cao et al. investigated PINNs for Poisson-like equations, without, however, extending their results to relevant electrical engineering problems [cao_physics-informed_2023]. Cai et al. proposed a TCAD forward simulation PINN-based architecture to solve a one-dimensional $p$-$n$ junction simulation. However, they did not discuss the generality of their method in higher dimensions or for inverse problems [cai_multi-order_2024]. Kim and Shin used a PINN-DeepONet hybrid solver to simulate the electrical characteristics of a one-dimensional device based on the TCAD training data [kim_novel_2023]. Lastly, Liu et al. developed a one-dimensional Poisson-Boltzmann solver, which addressed parameter retrieval problems [liu_asymptotic-preserving_2025]. As a result, there remains a critical need for a forward and inverse (unified) PINN framework that can address relevant devices and coupled simulations with inverse design capabilities. Such a framework should match the precision and reliability of state-of-the-art industrial solvers while demonstrating greater flexibility and scalability in integrating different boundary conditions, generation-recombination rates, and inversely retrieving relevant device parameters. Moreover, an innovative physics-informed deep learning framework should also demonstrate a clear computational advantage in solving relevant non-linear inverse problems beyond the reach of current numerical methods [noauthor_machine_2025]. As recently pointed out, PINNs should not be employed to solve relatively simple forward problems against a weak baseline without demonstrating a clear computational advantage over traditional solvers [mcgreivy_weak_2024]. Specifically, we argue that the crucial advantage of PINNs compared to traditional solvers stems from their inherent ability to successfully tackle inverse design problems where physics-constrained solutions, representing targeted functionalities or device configurations, are efficiently recovered even in the case of stiff and ill-posed problems [lu_physics-informed_2021, riganti_multiscale_2025]. Furthermore, there have been increasing efforts to extend the applications of physics-based surrogate models to benchmark performance applications, such as in the automotive aerodynamics field [tangsali_benchmarking_2025], stressing the importance of physics-informed deep learning for industrial applications.
@@
In addition, fields such as fluid dynamics and electromagnetics have experienced substantial advances in physics-informed surrogate modeling over the past five years, despite the availability of established forward Navier-Stokes or electrodynamics numerical solvers [comsol_sftw, chen_physics-informed_2020, eivazi_physics-informed_2022, chen_physics-informed_2022, riganti_multiscale_2025, mao_physics-informed_2020, raissi_hidden_2020]. In contrast, the comparatively limited progress of PINNs in semiconductor device modeling originates from the intrinsic challenges associated with the highly non-linear nature of the drift–diffusion equations, whose solutions present large gradients and span extremely large dynamic ranges. In our view, these considerations underscore the need for fundamentally new network architectures and training methodologies.
@@
While DDNet meets the essential benchmarks for comparison with current state-of-the-art modeling tools, it is not intended as a tout court replacement of established numerical methods. Instead, our comparisons should be regarded as seminal steps toward the development of PINN-driven semiconductor device modeling techniques enabling device discovery at minimal computational costs. Ultimately, our results unleash the potential of DDNet as a flexible and generalizable framework for microelectronics research, inverse design, and device discovery.
@@ 2 Results
2.1 The DDNet framework for semiconductor modeling
2.1.1 The Poisson Drift-Diffusion Equations
DDNet solves the general stationary drift-diffusion system for $\phi,n,$ and $p$:
@@ 2.1.1 The Poisson Drift-Diffusion Equations
The Poisson-drift-diffusion equations describe charge transport in semiconductors by coupling the Poisson equation for the electrostatic potential $\phi$ to the continuity equations for the electrons ($n$) and holes ($p$) concentrations. To solve Eq. 1 we must specify the device permittivity $\epsilon$ and doping profile $N_{d}^{+}-N_{a}^{-}. CC$ is the distribution of dopants within a semiconductor material, and it determines the electrical properties of the device. $\boldsymbol{J}_{n},\boldsymbol{J}_{p}$ denote the electron and hole current densities, respectively. Their divergences are linked to the generation–recombination term $R$ that represents the net rate of electron–hole pair creation or annihilation per unit volume. In the simulations that follow, we include the Shockley-Read-Hall recombination term, but DDNet has the flexibility to incorporate other mechanisms such as Auger or radiative effects. Charge transport is dominated by two phenomena. Charges driven by the electric field give rise to the drift currents while gradients in charge density originate diffusion currents, with mobilities and diffusivities linked via the Einstein relation [1985Markowich, sze_physics_2021]. Solving the drift-diffusion model in Eq. 1 requires the specification of appropriate boundary conditions, which could be Dirichlet, Neumann, or mixed depending on specific device structures. DDNet is designed to simultaneously learn the scaled functions $\hat{\phi}$, $\hat{n}$, and $\hat{p}$ [1985Markowich] (see the “Methods” section).
2.1.2 The DDNet architecture
[…]
Each loss term $\mathcal{L}_{t}$ is computed using a mean squared error formula to minimize the residual error on $\mathcal{N}_{i}$ collocation or boundary points using the ADAM optimizer [kingma_adam_2017]:
[…]
After the optimization step is completed, the network parameters $\tilde{\theta}$ are updated using backpropagation, and the process is repeated until a threshold error is achieved, or the training routine has reached a set number of epochs. For all the examples that follow, the loss functions of DDNet achieve values lower than $10^{-4}$, without the need for device- or geometry-specific fine-tuning, as shown in Table S2 and Figure S4 of the Supplementary Information.
@@ 2.1.2 The DDNet architecture
We present here prototypical examples of semiconductor device configurations in order to demonstrate the capabilities of the developed DDNet framework and we compare its predictions with those obtained from state-of-the-art commercial solvers. The results were compared using the relative $L_{1}$ error defined in Section 2 of the Supplementary Information. The following results comprise both forward and inverse design problems, including simulations of 1D and 2D homogeneous silicon junctions under varying bias conditions.
@@ 2.2 One-dimensional $p$-$n$ junction
We first simulate a silicon 1D $p$-$n$ junction at equilibrium using an abrupt doping profile. Despite having an analytical solution [1985Markowich, sze_physics_2021], this case is computationally challenging for traditional PINNs due to the presence of a discontinuity in $C$. As shown in Fig. 2a, the junction forms at the interface between $n$-type and $p$-type regions. At thermal equilibrium, the distribution of carriers creates a depletion region and establishes the built-in potential $\phi_{bi}$. Dirichlet Ohmic boundary conditions are applied for all fields to ensure uniqueness of the solution. Fig. 2b compares the DDNet prediction of $\phi$ with TCAD Sentaurus results, showing excellent agreement in logarithmic scale. The carrier density profiles are shown in Fig. 2c–d, achieving relative $L_{1}$ errors below 2%. We also simulated the same junction under forward and reverse bias, and included the results in the Supplementary Information Figure S3. DDNet well-captures the evolution of potential and carrier profiles for various applied voltages. In particular, DDNet captures the expected shift of the depletion region. Additionally, we simulate and include in the Supplementary Information Figure S3 an asymmetric ($p^{+}$-$n$) junction simulation, where the depletion region shifts toward the less-doped side.
@@ 2.3 Two-dimensional $p$-$n$ junction
To demonstrate scalability to higher spatial dimensions, we simulate a 2D $p$-$n$ junction with the geometry shown in Fig. 3a and doping profile in Fig. 3b. The boundary conditions are schematized in panel c, where Dirichlet conditions are enforced on the device contacts and Neumann insulating conditions elsewhere, modeling zero current outflow. The DDNet predictions for $\phi$, $n$, and $p$ are shown in panels e, h, and k, respectively, and compared to TCAD Sentaurus results in panels d, g, and j. The 3D surface plots (panels f, i, l) confirm accurate reconstructions, with the average relative $L_{1}$ errors with TCAD Sentaurus’s solutions falling below 5%. This example demonstrates robust scalability to higher dimensions, highlighting the expressivity of DDNet. In fact, the architecture, collocation points, and loss functions were left unchanged for the solution of the 2D problem.
@@
Figure 4: Inverse design capabilities of DDNet. a, Modified DDNet architecture to tackle inverse design and parametric learning tasks. The input space has been expanded to accept device parameters $P_{1},\dots,P_{n}$, and the output space of the $p$-Net and $n$-Net now enables the retrieval of the doping profile $\hat{C}=\hat{N}_{d}^{+}+\hat{N}_{a}^{-}$. b, Example of parametric learning for a pinpin junction, where DDNet extrapolates the maximal electric field behavior as a function of the intrinsically doped region width after training on 4 widths. c, Example of inverse design of a doping profile to achieve the maximal user-defined target performance from only three data points. d, Example of device discovery, where a pnppnp BJT-like structure is retrieved to match the user’s performance criteria on the electric field and carrier concentrations.
2.4 Inverse Design
2.4.1 DDNet for inverse problems and parametric learning